OpenStudy (anonymous):

what is comparing fractions by converting them into equivalent? sorry but I am living in Greece and I am American and sometimes I can't understand the meaning

OpenStudy (bahrom7893):

you have to convert them into the same form and then compare the numerators and denominators

OpenStudy (anonymous):

so what if it was 13/50 with 21/100

OpenStudy (bahrom7893):

so convert the bottom into the same number, say 100 right: 13/50 = 13*2/50*2 (im multiplying by 2/2=1, so im not changing the problem)

OpenStudy (bahrom7893):

so: 26/100 and 21/100 compare the numerators: 26>21, so 26/100 > 21/100

OpenStudy (anonymous):

Thank you that was easy just to times it by to get the same but the next one is more difficult 8/11 and 3/6